# How to Train an Oscillator Ising Machine using Equilibrium Propagation

### v2

@@ Abstract
We show that Oscillator Ising Machines (OIMs) are prime candidates for use as neuromorphic machine learning processors with Equilibrium Propagation (EP) based on-chip learning. The inherent energy gradient descent dynamics of OIMs, combined with their standard CMOS implementation using existing fabrication processes, provide a natural substrate for EP learning. Our simulations confirm that OIMs satisfy the gradient-descending update property necessary for a scalable Equilibrium Propagation implementation and achieve $\sim$97.2±0.1test accuracy on MNIST and $\sim$88.0±0.1on Fashion-MNIST without requiring any significant hardware modifications. Importantly, OIMs maintain robust performance under realistic hardware constraints, including 10-bit parameter quantization, 4-bit phase measurement precision, and moderate phase noise that can potentially be beneficial with parameter optimization. These results establish OIMs as a promising platform for fast and energy-efficient neuromorphic computing, potentially enabling energy-based learning algorithms that have been previously constrained by computational limitations.
@@ III Equilibrium Propagation
This total energy decomposes as $F=E+\beta\ell$, where $E(x,s,\{\theta\})$ is the free system’s energy function, $\ell(y,\hat{y})$ is a loss function comparing outputs $y$ to targets $\hat{y}$, and $\beta$ is the nudging factor that biases dynamics toward configurations that minimize loss.
@@
^$\hat{\nabla}^{\rm EP}(\beta)=-\frac{1}{2\beta}\left(\frac{\partial F}{\partial\theta}(x,s^{\beta}_{*},\{\theta\})-\frac{\partial F}{\partial\theta}(x,s_{*}^{-\beta},\{\theta\})\right).$ (4)
@@
$\lim_{\beta\to 0}\hat{\nabla}^{\rm EP}(\beta)=-\frac{\partial\ell}{\partial\theta}(y_{*},\hat{y}).$ (5)
@@ IV Implementation on Oscillator Ising Machines
We implement neural network components through specific energy terms in the OIM: biases via $-b_{i}\cos(\phi_{i})$ terms, and hidden-to-output weights via $-w^{(h,y)}_{ij}\cos(\phi_{i}^{(h)}-\phi_{j}^{(y)})$ terms. Using the mapping $s_{i}=\cos(\phi_{i})$ to transform phases into values in $[-1,1]$, we implement input-to-hidden weights through $-w^{(x,h)}_{ij}x_{i}\cos(\phi^{(h)}_{j})$ terms.
@@
$\frac{1}{2}\big{(}\cos(\phi^{(y)}_{i})-\hat{y}_{i}\big{)}^{2}$ $=\frac{1}{4}\cos(2\phi_{i}^{(y)})-\hat{y}_{i}\cos(\phi^{(y)}_{i})+\text{const.}$. (7)
@@ V Results
For practical hardware implementation, robustness to physical limitations is essential. Fig. 3 examines the effects of limited phase measurement precision for EP gradient computation. Here, $n$-bit phase measurement refers to quantizing converged phase values to $2^{n}$ evenly-spaced values between 0 and $2\pi$: $\phi_{*}$ before initializing the nudged phases, and $\phi_{*}^{\pm\beta}$ before applying the EP update equations (12)–(14). Remarkably, the network maintains full performance down to 4-bit phase measurement (89.8±1.5%), with significant degradation only at 2-bit precision.
@@
Similarly, Fig. 4 shows strong performance with 10-bit parameter precision (89.4±1.5%), with significant degradation only at 8-bit precision. Here, $n$-bit precision refers to the rounding of coupling, bias, and synchronization parameters to $2^{n}$ evenly-spaced values within their typical training ranges (determined from full-precision runs) before each dynamics phase. All physical dynamics evolve with these quantized parameters, and gradient updates are computed from the resulting converged phase values, then quantized before application.
@@
Fig. 5 demonstrates noise tolerance through proper parameter tuning. We model phase noise by adding Gaussian noise $\xi\mathcal{N}(0,1)$ to the phase dynamics equation (1) at each timestep. The top panel shows performance degradation with fixed hyperparameters ($\beta=0.05$) optimized for noise-free conditions, with accuracy declining from 90.8±0.9% to 41.0±6.2% as noise increases. However, the bottom panel reveals that when $\beta$ is optimized for each noise level, performance remains robust with no significant degradation and may even improve slightly. The network achieves 92.0±0.3% at $\xi=0.2$ with $\beta=0.3$, maintaining performance even at $\xi=0.3$ (unlikely in well-designed implementations, as shown in the right inset). We find that robust performance is maintained when $\beta\gtrsim\xi/2$ (left inset), though theoretical EP guarantees strictly apply in the small $\beta$ limit.

### v5

@@ Abstract
We show that Oscillator Ising Machines (OIMs) are prime candidates for use as neuromorphic machine learning processors with Equilibrium Propagation (EP) based on-chip learning. The inherent energy gradient descent dynamics of OIMs, combined with their standard CMOS implementation using existing fabrication processes, provide a natural substrate for EP learning. Our simulations confirm that OIMs satisfy the gradient-descending update property necessary for a scalable Equilibrium Propagation implementation and achieve $\sim$97.2±0.1% test accuracy on MNIST and $\sim$88.0±0.1% on Fashion-MNIST without requiring any significant hardware modifications. Importantly, OIMs maintain robust performance under realistic hardware constraints, including 10-bit parameter quantization, 4-bit phase measurement precision, and moderate phase noise that can potentially be beneficial with parameter optimization. These results establish OIMs as a promising platform for fast and energy-efficient neuromorphic computing, potentially enabling energy-based learning algorithms that have been previously constrained by computational limitations.
@@ III Equilibrium Propagation
[…]
@@
$\hat{\nabla}^{\rm EP}(\beta)=-\frac{1}{2\beta}\left(\frac{\partial F}{\partial\theta}(x,s^{\beta}_{*},\{\theta\})-\frac{\partial F}{\partial\theta}(x,s_{*}^{-\beta},\{\theta\})\right).$ (4)
@@
[…]
@@ IV Implementation on Oscillator Ising Machines
We implement neural network components through specific energy terms in the OIM: biases via $b_{i}\cos(\phi_{i})$ terms, and hidden-to-output weights via $w^{(h,y)}_{ij}\cos(\phi_{i}^{(h)}-\phi_{j}^{(y)})$ terms. Using the mapping $s_{i}=\cos(\phi_{i})$ to transform phases into values in [1,1][-1,1], we implement input-to-hidden weights through $w^{(x,h)}_{ij}x_{i}\cos(\phi^{(h)}_{j})$ terms.
@@
$\frac{1}{2}\big{(}\cos(\phi^{(y)}_{i})-\hat{y}_{i}\big{)}^{2}$ $=\frac{1}{4}\cos(2\phi_{i}^{(y)})-\hat{y}_{i}\cos(\phi^{(y)}_{i})+\text{const.}$ (7)
@@ V Results
For practical hardware implementation, robustness to physical limitations is essential. Fig. 3 examines the effects of limited phase measurement precision for EP gradient computation. Here, $n$-bit phase measurement refers to quantizing converged phase values to $2^{n}$ evenly-spaced values between 0 and $2\pi$: $\phi_{*}$ before initializing the nudged phases, and $\phi_{*}^{\pm\beta}$ before applying the EP update equations (12)–(14). Remarkably, the network maintains full performance down to 4-bit phase measurement (89.8±1.5% on MNIST/100), with significant degradation only at 2-bit precision.
@@
Similarly, Fig. 4 shows strong performance with 10-bit parameter precision (89.4±1.5% on MNIST/100), with significant degradation only at 8-bit precision. Here, $n$-bit precision refers to the rounding of coupling, bias, and synchronization parameters to $2^{n}$ evenly-spaced values within their typical training ranges (determined from full-precision runs) before each dynamics phase. All physical dynamics evolve with these quantized parameters, and gradient updates are computed from the resulting converged phase values, then quantized before application.
@@
Fig. 5 demonstrates noise tolerance through proper parameter tuning. We model phase noise by adding Gaussian noise $\xi\mathcal{N}(0,1)$ to the phase dynamics equation (1) at each timestep. The top panel shows performance degradation with fixed hyperparameters ($\beta=0.05$) optimized for noise-free conditions, with accuracy declining from 90.8±0.9% to 41.0±6.2% on MNIST/100 as noise increases. However, the bottom panel reveals that when $\beta$ is optimized for each noise level, performance remains robust with no significant degradation and may even improve slightly. The network achieves 92.0±0.3% at $\xi=0.2$ with $\beta=0.3$, maintaining performance even at $\xi=0.3$ (unlikely in well-designed implementations, as shown in the right inset). We find that robust performance is maintained when $\beta\gtrsim\xi/2$ (left inset), though theoretical EP guarantees strictly apply in the small $\beta$ limit.